The difference between two numbers is 121. 6 times the smaller is equal to 4 more than the larger. what are the numbers?
Answer by mananth(16949)

1 x + -1 y = 121 .............1
-1 x + 6 y = 4 .............2
Eliminate y
multiply (1)by 6
Multiply (2) by 1
6 x + -6 y = 726
-1 x + 6 y = 4
Add the two equations
5 x = 730
/ 5
x=146 =146.00
plug value of x in (1)
x-y = 121
146-y=121
-1 y = 121 + -146
-1 y = -25
y = 25